# Detection of disk-jet co-precession in a tidal disruption event

### v1

@@ Acknowledgments
We thank Pu Du for coordinating the Lijiang 2.4m Telescope observations, which provided two spectra taken on 2024-02-24. We also thank Ning Jiang, Youjun Lu, Song Wang, and Jianmin Wang for the useful discussion.
@@
This research was supported by the National Natural Science Foundation of China (NSFC) under grant numbers 11988101, 11933004, 12173103, and 12261141691; by the National Key Research and Development Program of China (NKRDPC) under grant numbers 2019YFA0405504 and 2019YFA0405000; by the Strategic Priority Program of the Chinese Academy of Sciences under grant numbers XDB41000000 and XDB0550203; by ANID, Millennium Science Initiative, ICN12_009; and by a research grant (VIL60862) from VILLUM FONDEN. SdP acknowledges support from ERC Advanced Grant 789410. P.C. acknowledges support via Research Council of Finland (grant 340613). SNZ acknowledges support from the National Natural Science Foundation of China (grant Nos. 12333007). MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. TP acknowledges the financial support from the Slovenian Research Agency (grants I0-0033, P1-0031, J1-8136, J1-2460 and Z1-1853). H.Z. acknowledges National Key R&D Program of China (grant Nos. 2023YFA1607804, 2022YFA1602902, and 2023YFA1608100), and National Natural Science Foundation of China (NSFC; grant Nos. 12120101003, 12373010, and 12233008). L.G. acknowledges financial support from AGAUR, CSIC, MCIN and AEI 10.13039/501100011033 under projects PID2023-151307NB-I00, PIE 20215AT016, CEX2020-001058-M, ILINK23001, COOPB2304, and 2021-SGR-01270. CPG acknowledges financial support from the Secretary of Universities and Research (Government of Catalonia) and by the Horizon 2020 Research and Innovation Programme of the European Union under the Marie Skłodowska-Curie and the Beatriu de Pinós 2021 BP 00168 programme, from the Spanish Ministerio de Ciencia e Innovación (MCIN) and the Agencia Estatal de Investigación (AEI) 10.13039/501100011033 under the PID2023-151307NB-I00 SNNEXT project, from Centro Superior de Investigaciones Científicas (CSIC) under the PIE project 20215AT016 and the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, and from the Departament de Recerca i Universitats de la Generalitat de Catalunya through the 2021-SGR-01270 grant. MPT acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S and from the Spanish grant PID2023-147883NB-C21, funded by MCIU/AEI/10.13039/501100011033, as well as support through ERDF/EU. T.-W.C. acknowledges the Yushan Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). FGX and LHW were supported in part by National SKA Program of China (grant Nos. 2020SKA0110100 and 2020SKA0110200), and by National Natural Science Foundation of China (NSFC; grant Nos. 12373017, 12192220 and 12192223). FO acknowledges support from MIUR, PRIN 2020 (grant 2020KB33TP) “Multimessenger astronomy in the Einstein Telescope Era (METE)” and from INAF-MINIGRANT (2023): “SeaTiDE - Searching for Tidal Disruption Events with ZTF: the Tidal Disruption Event population in the era of wide field surveys”.
@@
YNW initiated the project, coordinated the acquisition of most used data, and contributed significantly to the text. ZKL and LHW analyzed X-ray and radio data, respectively, and contributed to the text. LWH initiated the modeling, estimated the spin, and contributed to text. SYW analyzed UV and optical data, modeled the host-galaxy SED, and contributed to the text. LJ, SNZ, JFL, BZ, FGX, DFB, XWC, HF, ELQ, XWS and DRP contributed to the physical interpretation. SDP modeled the radio SED and contributed to the text. RDB analyzed VLBA data, helped with VLA data and contributed to the text. YH contributed to optical observations, together with YWZ, estimated the black hole mass, and contributed to the text. AYY, RRC, LY, and ALW analyzed radio data. HZW contributed to estimate the black hole mass. PC contributed to NTT spectral analysis and the text. DW helped with radio data analysis. ZHY, JW, YLQ and JYW contributed to SVOM data and the text. WXL, HZ and WJG contributed to DESI data. GL, JPA, MN, TW, TEM, CPG, MG, CI, LM, FO, TP, LG, MPT and TWC are ePESSTO+ builders. DA contributed to data analysis. All co-authors contributed with comments to the text.
@@
All NICER and Swift data are available from NASA’s High Energy Astrophysics Science Archive Research Center (https://heasarc.gsfc.nasa.gov/). The XMM-Newton data are currently under the proprietary and will be publicly available in February, 2025. The data used to generate Figs. 1, 2, 4, and 5, along with the Python codes for the Lomb-Scargle periodogram and cross-correlation function, are available on Zenodo (https://doi.org/10.5281/zenodo.14195067).
@@
Sections S1 to S9
Figures S1 to S8
Table S1
@@
Sections S1 to S9
Figures S1 to S8
Table S1
Supplementary
@@ 1 Observations and Data Analysis
AT2020afhd was well detected by both ground- and space-based facilities, with observations spanning from radio to soft X-rays (see the target’s localization in Extended Data Figure, Fig. S1). Around late November 2024, AT2020afhd exhibited another re-brightening across multiwavelengths at the time of writing. In this study, we focused exclusively on the period between January 1 and November 26, 2024, before the onset of the second re-brightening.
[…]
@@ 1.1.1 Ultraviolet Optical Telescope (UVOT)/Neil Gehrels Swift Observatory (Swift)
Swift (?) performed 66 target-of-opportunity (ToO) observations on this target from January 31 to November 22, 2024 (PI: Y. Wang, E. Hammerstein, N. Jiang, and S. Huang). When multiple snapshots were included in the observations of each individual filter, the uvotimsum task was used to combine all exposures. We used an 8” aperture centered on the target position as the source, and a nearby 40” source-free region as the background. The target was observed with all or some of the six filters–UVW2 (central wavelength, 1928 Å), UVM2 (2246 Å), UVW1 (2600 Å), U (3465 Å), B (4392 Å) and V (5468 Å) –across different observations. The UVOT magnitudes have been corrected for Galactic extinction and host galaxy contributions (see the host galaxy properties in the Supplementary).
@@ 1.1.2 Asteroid Terrestrial-impact Last Alert System (ATLAS)
The ATLAS telescopes conducted observations by scanning the visible sky with a cadence of one to two nights, using broadband filters as part of the standard asteroid search program (?, ?). Specifically, two filters were employed: the c-band (4200–6500 Å) and o-band (5600–8200 Å). We obtained the difference photometry data of the source from the ATLAS forced photometry server (?) and selected observations with MJD >> 60300 and a signal-to-noise ratio (SNR) >> 5 to construct the lightcurve within the studied period. To enhance the accuracy of flux measurements, intra-night observations were combined into several epochs using a weighted mean, with weights assigned as the inverse variances $1/\sigma^{2}$ of each individual observation. The uncertainty for each epoch was calculated by the following formula: σ_epoch^2 = 1∑i=1N(1σi2) + 1N-1∑_i=1^N(m_i - m_epoch)^2 We applied iterative 3-sigma clipping within each epoch to remove outliers from intra-night observations. Due to its superior sampling, only the o-band data were included in this work. Finally, we selected epochs with $\frac{m_{\rm epoch}}{\sigma_{\rm epoch}}>3$ and corrected those magnitudes for Galactic extinction.
@@ 1.1.4 New Technology Telescope (NTT)
Optical spectra of AT2020afhd were extensively collected using BFOSC with a new Grism G4, mounted on the 2.16-meter telescope at the Xinglong Observatory in Hebei Province, operated by the National Astronomical Observatories of the Chinese Academy of Sciences. Between February 9 and March 20, 2024, a total of 11 optical spectra covering the wavelength range of 385075003850-7500 Å  were obtained. The typical exposure times ranged from 1800 to 2400 seconds, depending on the specific weather conditions. Most spectra were observed using a slit width of $1.8^{\prime\prime}$, corresponding to a resolution of approximately 20~20 Å. However, a few spectra were taken with a slit width of $2.3^{\prime\prime}$, yielding a resolution of $25$ Å, due to poor seeing conditions. After bias subtraction and flat-field correction, the spectra were reduced using a specialized Python package specifically designed for the BFOSC instrument on the 2.16-m telescope.
@@ 1.1.7 Dark Energy Spectroscopic Instrument (DESI)
DESI is amounted on the 4-meter Mayall Telescope, located at the Kitt Peak National Observatory in Arizona, USA (?, ?). On February 19, 2022, DESI successfully obtained one spectrum of AT2020afhd. This spectrum was selected from the DESI Bright Galaxy Survey (?), with target sources coming from the DESI Legacy Imaging Survey (?, ?). The spectral range spans from 3600 Å  to 9800 Å, with an exposure time of $780$ s and a spectral resolution of R=20005000R=2000-5000. The observed data were processed using the DESI spectroscopic data processing pipeline (?).
@@ 1.2.1 X-Ray Telescope (XRT)/Swift
XRT observations were performed in photon-counting mode, with a total exposure time of $\sim 132.6$ ks. The XRT spectra were downloaded directly from the UK Swift Science Data Centre (?, ?). To constrain the continuum, only observations with more than 8 counts were included in the analysis. The spectra were grouped using the FTOOL ftgrouppha with grouptype=optmin and groupscale=3. Consequently, Cash statistics ( (?)) was applied for spectral fitting. As the source is background-dominated above 2 keV, the XRT spectra were fitted in the 0.3–2 keV energy band only.
@@ 1.2.2 Neutron Star Interior Composition Explorer (NICER)
NICER observed AT2020afhd from February 4, to November 24, 2024, with a total exposure time of 156.9 ks (PI: Y. Wang, Y. Yao, and T. Wevers). Raw data were obtained from the HEASARC public archive and processed using the NICER Data Analysis Software (version v2023-08-22_V011a). Calibration (version xti20221001) and screening of the raw data were performed using nicerl2 with thresh_range=“–3.0–38”. GTIs with count rates exceeding 1 count per second in the 8–12 keV band were excluded to prevent contamination from background flares. For observations with a single GTI and a gap of less than 0.1 days before the next observation, we merged the two using niobsmerge to improve the SNR. Spectra were generated with nicerl3-spect, and background contributions were estimated using the SCORPEON model.***https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-overview/. The spectra were binned with ftgrouppha, using grouptype=optmin and groupscale=3, ensuring a minimum of 3 counts per bin. For spectral fitting, the energy range of 0.3–15 keV was used, and the con_norm parameter was allowed to vary to accurately estimate the background components.

### v2

@@ Acknowledgments
We thank Pu Du for coordinating the Lijiang 2.4m Telescope observations, which provided two spectra taken on 2024-02-24. We also thank Ning Jiang, Wenda Zhang, Youjun Lu, Song Wang, and Jianmin Wang for the helpful discussions, and Wu Jiang for his help with the VLBA data analysis.
@@
This research was supported by the National Natural Science Foundation of China (NSFC) under grant numbers 12588202, 12173103, 12261141691, 12473012 and 12203041; by the New Cornerstone Science Foundation through the New Cornerstone Investigator Program and the XPLORER PRIZE; by the Strategic Priority Program of the Chinese Academy of Sciences under grant numbers XDB41000000 and XDB0550203; by ANID, Millennium Science Initiative, ICN12_009; and by a research grant (VIL60862) from VILLUM FONDEN. SdP acknowledges support from ERC Advanced Grant 789410. P.C. acknowledges support via Research Council of Finland (grant 340613). SNZ acknowledges support from the National Natural Science Foundation of China (grant Nos. 12333007). MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. TP acknowledges the financial support from the Slovenian Research Agency (grants I0-0033, P1-0031, J1-8136, J1-2460 and Z1-1853). H.Z. acknowledges National Key R&D Program of China (grant Nos. 2023YFA1607804, 2022YFA1602902, and 2023YFA1608100), and National Natural Science Foundation of China (NSFC; grant Nos. 12120101003, 12373010, and 12233008). L.G. acknowledges financial support from AGAUR, CSIC, MCIN and AEI 10.13039/501100011033 under projects PID2023-151307NB-I00, PIE 20215AT016, CEX2020-001058-M, ILINK23001, COOPB2304, and 2021-SGR-01270. CPG acknowledges financial support from the Secretary of Universities and Research (Government of Catalonia) and by the Horizon 2020 Research and Innovation Programme of the European Union under the Marie Skłodowska-Curie and the Beatriu de Pinós 2021 BP 00168 programme, from the Spanish Ministerio de Ciencia e Innovación (MCIN) and the Agencia Estatal de Investigación (AEI) 10.13039/501100011033 under the PID2023-151307NB-I00 SNNEXT project, from Centro Superior de Investigaciones Científicas (CSIC) under the PIE project 20215AT016 and the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, and from the Departament de Recerca i Universitats de la Generalitat de Catalunya through the 2021-SGR-01270 grant. MPT acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S and from the Spanish grant PID2023-147883NB-C21, funded by MCIU/AEI/10.13039/501100011033, as well as support through ERDF/EU. T.-W.C. acknowledges the Yushan Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). FGX and LHW were supported in part by National SKA Program of China (grant Nos. 2020SKA0110100 and 2020SKA0110200), and by National Natural Science Foundation of China (NSFC; grant Nos. 12373017, 12192220 and 12192223). FO acknowledges support from MIUR, PRIN 2020 (grant 2020KB33TP) “Multimessenger astronomy in the Einstein Telescope Era (METE)” and from INAF-MINIGRANT (2023): “SeaTiDE - Searching for Tidal Disruption Events with ZTF: the Tidal Disruption Event population in the era of wide field surveys”.
@@
Conceptualization: Y.N.W., W.H.L., S.N.Z., J.F.L., S.D.P., R.D.B., A.Y.Y., X.W.C., E.L.Q., X.W.S., F.O., and D.A. Methodology: Y.N.W., Z.K.L., W.H.L., S.N.Z., L.J., J.F.L., B.Z., F.G.X., D.F.B., D.R.P., S.D.P., A.Y.Y., and D.A. Software: Y.N.W., Z.K.L., L.H.W., W.H.L., S.D.P., Y.H., and A.Y.Y. Validation: Y.N.W., Z.K.L., L.H.W., W.H.L., Y.H., A.Y.Y., P.C., Y.W.Z., W.X.L., H.Z.W., T.E.M., and J.Y.W. Formal analysis: Y.N.W., Z.K.L., L.H.W., W.H.L., L.J., F.G.X., Y.H., A.Y.Y., R.R.C., Y.W.Z., X.W.C., H.Z.W., D.W., and Z.H.Y. Investigation: Y.N.W., Z.K.L., W.H.L., S.D.P., R.D.B., L.Y., Z.H.Y., F.G.X., D.F.B., X.W.C., H.F., Y.H., A.Y.Y., A.L.W., P.C., D.W., W.X.L., M.N., G.L., C.P.G., T.P., T.W., J.Y.W., W.J.G., L.G., and T.W.C. Resources: Y.N.W., L.H.W., Y.H., A.Y.Y., J.W., H.Z., W.J.G., G.L., J.P.A., T.E.M., Y.L.Q., C.I., T.P., L.G., and T.W.C. Data curation: Y.N.W., Z.K.L., L.H.W., S.Y.W., A.Y.Y., L.Y., A.L.W., Y.W.Z., Z.H.Y., J.W., G.L., W.J.G., C.I., J.P.A., M.G., and T.P. Writing—original draft: Y.N.W., Z.K.L., L.H.W., W.H.L., L.J., J.F.L., S.D.P., R.D.B., D.F.B., Z.H.Y., and H.Z. Writing—reviewing and editing: Y.N.W., Z.K.L., L.H.W., W.H.L., L.J., S.N.Z., J.F.L., F.G.X., D.F.B., H.F., S.D.P., X.W.S., R.D.B., Y.H., A.Y.Y., P.C., D.W., G.L., J.P.A., M.N., L.M., T.E.M., C.P.G., T.W., M.G., F.O., D.A., L.G., M.P.T., and T.W.C. Visualization: Y.N.W., Z.K.L., L.H.W., W.H.L., S.D.P., R.D.B., J.Y.W., and P.C. Supervision: Y.N.W., W.H.L., and J.F.L. Project administration: Y.N.W., J.F.L., and J.P.A. Funding acquisition: Y.N.W., W.H.L., J.F.L., and C.I.
@@
The data used to generate Figs. 1, 2, 3, and 5, along with the Python codes for the Lomb-Scargle periodogram and cross-correlation function, are available on Zenodo (https://doi.org/10.5281/zenodo.17649346).
@@
Sections S1 to S6
Figures S1 to S5
[…]
Figure 2: Temporal evolution of spectral parameters. (A) The diskbb temperature derived from X-ray spectra. The luminosity evolution is depicted with grey symbols to illustrate its correlation with temperature; the two parameters generally co-evolve, with the temperature reaching its peak alongside luminosity during the first 300 days. (B) The in-band spectral index ($F\propto\nu^{\alpha}$) derived from the VLA (4–8 GHz) and ATCA (4–10 GHz) data. The stars indicate the VLA observations used for the radio SED modeling (Supplementary 3). (C) The peak frequency, $\nu_{\rm p}$, and the peak flux, $F_{\rm p}$, derived from the SED modeling.
Figure 3: Timing analysis of the X-ray and radio data. (A) Lomb-Scargle periodogram of the X-ray lightcurve. The calculation includes data collected between August 3 and October 21, 2025, during which the X-ray quasi-periodic variations were clearly observed. (B) Cross-correlation function between the X-ray and radio data. The data used in this calculation are indicated by the gray-shaded region in Figs. 1A and B. The histogram represents the distribution obtained from bootstrap simulations, with the red dashed line marking the median of the distribution at 19.0-19.0 days. (C) Folded X-ray and radio lightcurves with a period of 19.6 days. The radio data were rebinned into 0.1-phase intervals using a weighted mean for clarity. The data used in this calculation are shown in Fig. 5B.
Figure 4: Comparison of TDEs with early intense radio detection. The luminosities of the two on-axis jetted TDEs, Swift J1644+57 and AT2022cmc, as well as AT2020afhd, have been rescaled for clarity. The AMI-LA data at 15.5 GHz for Swift J1644+57, ASSASN–14li, AT2019azh and AT2022cmc were adapted from (?), (?), (?), and (?), respectively.
Figure 5: The disk-jet precession model. (A) Schematics of the proposed disk-jet precession model. $\theta_{\rm obs}$ and $\theta_{\rm i}$ represent the viewing angle of the system and the disk/jet precession angle around the black hole axis, respectively. The left and right plots correspond to the phases of the X-ray and radio variations when the luminosity is relatively low and high, respectively. (B) Comparison of the disk-jet precession model (the lower and upper black curves) with X-ray (0.1–2 keV) and radio (5–6 GHz) observations. In the presented model, we adopted a BH mass of $M_{\bullet}=10^{6.7}M_{\odot}$, a scaleheight ratio of $H/R\sim 1$, and an outer disk radius equal to the circularization radius of the debris. As a result, we determined an inclination angle of $\theta_{\rm obs}\sim 37.8-38.9^{\circ}$, a disk/jet precession angle of $\theta_{\rm i}\sim 14-15^{\circ}$, and a Doppler factor of $\Gamma\sim 1.2-1.6$.
Figure 6: Estimation of system parameters. (A) Disk precession period vs BH spin when adopting various BH masses and surface density profile of the accretion disk. The gray and the blue shades correspond to a BH mass of $10^{6.7\pm 0.5}~M_{\odot}$ and a period of $19.6\pm 1.5$ days. The overlap of the two shades constrains the BH spin to ranges of $0.46<a_{\bullet}<-0.14$ and $0.11<a_{\bullet}<0.35$. (B) Contour plots showing the best-fitting parameters for our proposed disk precession model.
[…]
Table 1: Summary of the spectroscopic observations
4d P60/SEDM - 2024-01-13T06:00:32 $2700$
43d FLOYDS-S - 2024-02-13T10:36:20 $1800$
[…]
Note. Spectra from P60 and FLOYDS were downloaded directly from the Transient Name Server (https://www.wis-tns.org/object/2020afhd). The epochs correspond to the times offset from MJD 60310. Spectra with a SNR below 10 are excluded from the table, as they are not utilized in the subsequent analysis.
Figure 8: White noise tests for LSP. (A) The CDF of the normalized LSP noise power from the observed data (blue points) compared to the expected CDF of an exponential distribution (red curve). (B) The PDF of the normalized LSP noise power from the observed data (blue histogram) compared to the expected PDF of an exponential distribution (red curve). (C) K-S test results from simulations. The orange histogram shows the distribution of K-S statistic values for simulated white noise, with the $1\sigma$ range shaded in grey. The red line represents the K-S statistic value for the observed LSP noise power. (D) A-D goodness-of-fit test results from simulations. The orange histogram shows the distribution of A-D statistic values for simulated white noise, with the $1\sigma$ range shaded in grey. The red line represents the A-D statistic value for the observed LSP noise power.
@@
Sections S1 to S6
Figures S1 to S5
Supplementary Materials
@@ 1.1.1 Ultraviolet Optical Telescope (UVOT)/Neil Gehrels Swift Observatory (Swift)
Swift (?) performed 66 target-of-opportunity (ToO) observations of AT2020afhd from January 31 to November 22, 2024 (PI: Y. Wang, E. Hammerstein, N. Jiang, and S. Huang). When multiple snapshots were included in the observations of each individual filter, the uvotimsum task was used to combine all exposures. We used an 8′′ aperture centered on the target position as the source, and a nearby 40′′ source-free region as the background. The target was observed with all or some of the six filters, i.e., UVW2 (central wavelength, 1928 Å), UVM2 (2246 Å), UVW1 (2600 Å), U (3465 Å), B (4392 Å) and V (5468 Å), across different observations. The UVOT magnitudes have been corrected for Galactic extinction and host galaxy contributions (see Supplementary 4).
@@ 1.1.2 Asteroid Terrestrial-impact Last Alert System (ATLAS)
The ATLAS telescopes conducted observations by scanning the visible sky with a cadence of one to two nights, using broadband filters as part of the standard asteroid search program (?, ?). Specifically, two filters were employed: the c-band (4200–6500 Å) and o-band (5600–8200 Å). We retrieved difference photometry data of the source from the ATLAS forced photometry server (?), selecting observations between MJD 60300 and 60640 with a signal-to-noise ratio (SNR) >> 5. To enhance the accuracy of flux measurements, intra-night observations were combined into several epochs using a weighted mean, with weights assigned as the inverse variances $1/\sigma^{2}$ of each individual observation. The uncertainty for each epoch was calculated by the following formula: σ_epoch^2 = 1∑i=1N(1σi2) + 1N-1∑_i=1^N(m_i - m_epoch)^2. We applied iterative 3-sigma clipping within each epoch to remove outliers from intra-night observations. Due to its superior sampling, only the o-band data were included in this work. Finally, we selected epochs with $\frac{m_{\rm epoch}}{\sigma_{\rm epoch}}>3$ and corrected those magnitudes for Galactic extinction.
@@ 1.1.4 New Technology Telescope (NTT)
Optical spectra of AT2020afhd were extensively collected using BFOSC with a new Grism G4, mounted on the 2.16-meter telescope at the Xinglong Observatory in Hebei Province, operated by the National Astronomical Observatories of the Chinese Academy of Sciences. Between February 9 and March 20, 2024, a total of 11 optical spectra covering the wavelength range of 3850–7500 Å  were obtained. The typical exposure times ranged from 1800 to 2400 seconds, depending on the specific weather conditions. Most spectra were observed using a slit width of $1.8^{\prime\prime}$, corresponding to a resolution of approximately 20~20 Å. However, a few spectra were taken with a slit width of $2.3^{\prime\prime}$, yielding a resolution of $25$ Å, due to poor seeing conditions. After bias subtraction and flat-field correction, the spectra were reduced using a specialized Python package specifically designed for the BFOSC instrument on the 2.16-m telescope.
@@ 1.1.7 Dark Energy Spectroscopic Instrument (DESI)
DESI is amounted on the 4-meter Mayall Telescope, located at the Kitt Peak National Observatory in Arizona, USA (?, ?). On February 19, 2022, DESI successfully obtained one spectrum of AT2020afhd. This spectrum was selected from the DESI Bright Galaxy Survey (?), with target sources coming from the DESI Legacy Imaging Survey (?, ?). The spectral range spans from 3600 Å  to 9800 Å, with an exposure time of $780$ s and a spectral resolution of $R$ = 2000–5000. The observed data were processed using the DESI spectroscopic data processing pipeline (?).
@@ 1.2.1 X-Ray Telescope (XRT)/Swift
XRT observations were performed in photon-counting mode, with a total exposure time of $\sim 132.6$ ks. The XRT spectra were downloaded directly from the UK Swift Science Data Centre (?, ?). To constrain the continuum, only observations with more than 8 counts were included in the analysis. The spectra were grouped using the FTOOL ftgrouppha with grouptype=optmin and groupscale=3. Consequently, Cash statistics (?) was applied for spectral fitting. As the source is background-dominated above 2 keV, the XRT spectra were fitted in the 0.3–2 keV energy band only.
@@ 1.2.2 Neutron Star Interior Composition Explorer (NICER)
NICER observed AT2020afhd from February 4, to November 24, 2024, with a total exposure time of 156.9 ks (PI: Y. Wang, Y. Yao, and T. Wevers). Raw data were obtained from the HEASARC public archive and processed using the NICER Data Analysis Software (version v2023-08-22_V011a). Calibration (version xti20221001) and screening of the raw data were performed using nicerl2 with thresh_range=“–3.0–38”. Good time intervals (GTIs) with count rates exceeding 1 count per second in the 8–12 keV band were excluded to prevent contamination from background flares. For observations with a single GTI and a gap of less than 0.1 days before the next observation, we merged the two using niobsmerge to improve the SNR. Spectra were generated with nicerl3-spect, and background contributions were estimated using the SCORPEON model (https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-overview/). The spectra were binned with ftgrouppha, using grouptype=optmin and groupscale=3, ensuring a minimum of 3 counts per bin. For spectral fitting, the energy range of 0.3–15 keV was used, and the con_norm parameter was allowed to vary to accurately estimate the background components.